OpenStudy (anonymous):

Multiply and simplify:

OpenStudy (anonymous):

\[\sqrt{3(7-}\sqrt{3)}\] is the problem.

OpenStudy (anonymous):

The answer choices are:

OpenStudy (anonymous):

A. \[\sqrt{21-}\sqrt{3}\] B. \[7\sqrt{3-}\sqrt{6}\] C. \[7\sqrt{3-3}\]

OpenStudy (lilg132):

A